Ali bicycles to a place 20km away. John's route to the same place is 4km longer but he rides 2km/h faster than Ali and gets their 5mins later. What are their speeds?
1
Expert's answer
2019-11-25T10:00:31-0500

Assume Ali's speed is "v" , then John's speed is "v + 2" , km/h.


Time spent: by Ali is "\cfrac{20}v" , by John is "\cfrac{20 + 4}{v + 2}" ., h.


Given that John spent 5 minutes (1/12 hour) more, we can make the equation


"\qquad \cfrac{20}{v} + \cfrac{1} {12} = \cfrac{20+4}{v + 2}"


To get rid of fractions we multiply the left and right sides of the equation by "12v(v+2)"


"\qquad 20 * 12 (v+2) + 1*v(v+2) = 24*12v"


"\qquad v^2 - 46v + 480 = 0"


"\qquad (v -16)(v - 30) = 0"


"\qquad \underline{v_1 = 16}, \space \underline{v_2 = 30}"


The problem has two solutions: 

1) Ali's speed is 16 km/h, John's speed is 18 km/h.

2) Ali's speed is 30, John's speed is 32.